OpenStudy (anonymous):

solve the following equation for x between 0 and 2pi 15 sin^2x+14 sinx -1=0

OpenStudy (anonymous):



OpenStudy (campbell_st):

factorise the quadratic (sin(x) +1)(15sin(x) -1) = 0 then sin(x) = -1 or 1/15 so sin(x) = 1/15 will have values in 1st and 2nd quadrants sin(x) = -1 is 270 degrees of 3pi/2 I'll let you figure out sin(x) = 1/15

OpenStudy (anonymous):

it's 0.0667

OpenStudy (anonymous):

but i have three answers 0.0667 , 3.0749 , 4.7124

OpenStudy (campbell_st):

thats radians... and only the 1st quadrant answer, you still need to find the 2nd quadrant answer because of the domain you stated

OpenStudy (anonymous):